OpenStudy (anonymous):

The amount of sales tax on a purchase varies directly as the price of the item varies. If a computer costs $400 before tax and the tax is $30 tax, which of the following equations is a correct step in solving for how much tax is needed for a computer that costs $500?

OpenStudy (anonymous):

are you saying that the tax percantage changes

OpenStudy (anonymous):

well my options are 400/30 = 500/y 400/30 = x/500 30/400 = 500/y 30x = 200,00 I think they're asking how to figure out how much tax

OpenStudy (anonymous):

400/30 = 500/y

OpenStudy (anonymous):

That's what I thought, thank you!!